# Beyond Worst-Case Subset Sum: An Adaptive, Structure-Aware Solver with Sub-$2^{n/2}$ Enumeration

### v2

@@ Abstract.
jesus.salas@gmail|microsoft.com
@@ 3.2. Unique Subset Sums
For clarity, we assume throughout that $S\subset\mathbb{Z}_{\geq 0}$. The methods generalize naturally to inputs containing 00 (introducing a trivial collision) or negative values (handled by symmetric memoization). In the worst case, storing all unique subset sums requires $O^{*}(2^{n/2})$ space, though this cost is typically much lower in collision-rich instances.
@@ 4. Prior Work
The Subset Sum problem has attracted extensive research due to its NP-completeness and practical significance. Early approaches, such as Bellman’s dynamic programming algorithm [16], offered pseudopolynomial-time solutions with time complexity $O(n\,t)$ (where $t$ is the target sum). Over time, several refined DP-based techniques have been proposed to mitigate this cost when $t$ remains large. For instance, Koiliaris and Xu [17] introduced a faster pseudopolynomial-time algorithm running in $\widetilde{O}(t\sqrt{n})$ by leveraging FFT-based convolutions for partial-sum computations. More recently, Bringmann [18] developed a near-linear pseudopolynomial algorithm for Subset Sum under specific parameter regimes, achieving further speedups in cases where the numeric range is not excessively large. Nevertheless, once $t$ grows beyond a moderate threshold, these DP-based methods become infeasible in practice because their complexity remains tied to the numerical size of the target.
@@ 5. Unique Subset Sums Enumerator
[…]
@@
[…]
@@
This strategy eliminates redundant enumeration and significantly reduces both computational overhead and memory usage, while ensuring that all unique subset sums—i.e., all elements of $\Sigma(S)$ (and hence the effective parameter $U$)—are captured.
@@ 5.6. Complement Trick for Half-Enumeration
Let $\ell$ be one half of the input, containing $n/2$ elements. Suppose the enumerator enumerates only subsets of size $\leq\lfloor n/4\rfloor$. Let $U_{\max}$ be the number of unique subset sums discovered. Then for any subset $S^{\prime}\subseteq\ell$ of size $>n/4$, its sum can be reconstructed via complement logic using previously enumerated subsets of size $\leq n/4$.
@@
Let $S^{\prime}\subseteq\ell$ be a subset of size $>n/4$. Then its complement $S^{*}=\ell\setminus S^{\prime}$ has size $<n/4$, since $\ell$ has size $n/2$. The enumerator processes all such small subsets and stores their sums. Let $\sigma^{*}=\sum_{x\in S^{*}}x$ be the recorded sum. Then:
@@ 6. Subset-Sum Solver with Double Meet-in-the-Middle Optimization
Let $\sigma(A):=\sum_{a\in A}a$ denote the sum of elements in a subset $A\subseteq S$, in contrast to $\Sigma(A)$, which refers to the full sumset of all subsets of $A$
@@ 6.1. Time Complexity Analysis
The enumeration algorithm maintains a dynamic frontier of reachable sums across both halves of the input. Each side’s unique sums are generated via the structure-aware subset sum enumerator described earlier, and combinations are performed only on compatible pairs. For every sum $s$ from the left half, the algorithm checks for a corresponding sum $t-s$ in the right half, or other lemma-induced matches.
@@ 7. Meet-in-the-Middle Speed-Up via Controlled Aliasing
This aliasing means that the distinct subsets $\{x_{0}\}$ and $\{x_{1}\}$ are both treated as contributing the same value to the total sum. Normally, the four subsets involving these two elements—$\emptyset$, $\{x_{0}\}$, $\{x_{1}\}$, and $\{x_{0},x_{1}\}$—can produce up to four different partial sums. Under our rule, the sums for $\{x_{0}\}$ and $\{x_{1}\}$ become identical, collapsing these states into fewer unique outcomes.
@@ 8.1. Overview of the Optimization
Figure 1 shows the number of $k$-subsets generated in each column expansion sub-process for an input instance with $n=48$. In this case, the process is executed $n/4=12$ times, yielding a total of 8 388 60783886078\,388\,6078 388 607 $k$-subsets per split, which corresponds to the expected $\frac{2^{n/2}}{2}$ in the worst case. The memoization structure holds only half of the unique subset sums, reflecting the true effective search space.
@@ 8.2. The Look-Ahead and Rescheduling Mechanism
[…]
@@ 9.1. Key Factors Influencing Instance Classification and Hardness
Let $n$ denote the number of elements in the set $S$, and let $w=\log_{2}(\max S)$ be the bit-length of the largest element. In addition to the classical measures of additive structure—such as the doubling constant $C$ and additive energy $E(S)$—we also consider the extended structural parameters introduced earlier: the linearity factor $\lambda$, the clustering factor $\gamma$, and the duplicate measure $\delta$. These parameters collectively capture the degree of additive redundancy in $S$ and offer a finer-grained understanding of how collisions among subset sums occur.
@@ 10.3. Duplicates Elements
We observe that the most expensive enumeration matches our baseline results, and that each additional duplicate reduces the number of effective unique subset sums by $\leavevmode\nobreak\ 25\%$ from the previous run. These sums are tracked using canonical bitmask representations, avoiding redundant credit for distinct but permutationally equivalent subsets.
@@ 12.1. Online Updates as a New Branch
A distinctive benefit of our anytime exploration is that it renders the search process highly governable. Since the algorithm operates in configurable discrete cycles—each producing a set of candidate branches that contribute to $U$—we can reassess the search state after every cycle. Once a cycle completes, the scheduler can reorganize the rescheduled branches by prioritizing those likely to yield new unique canonical subset sums, pruning those that are redundant (i.e., those whose sums already exist in $\Sigma(S)$), and even deliberately encouraging collisions to block redundant exploration in unpromising regions of the search space. This adaptive prioritization not only guarantees the delivery of partial solutions at any interruption point but also enables the solver to be controlled and fine-tuned in real time, enhancing overall efficiency.
@@ Appendix A Unique Subset Sums Enumerator Algorithm
[…]
2:sum $\leftarrow 0$triangleright Current subset sum
3:elements[n] $\leftarrow\{0,0,\ldots\}$triangleright Inclusion bitmask
@@
6:$S\leftarrow\{\ldots\}$
7:$n\leftarrow\text{len}(S)$
8:$MemoizedSums\leftarrow\{\}$
9:$CanonicalSubsets\leftarrow\{\}$
10:$Input\leftarrow\{\texttt{kPerm}(0,\{\})\}$
@@
25:                else if $expanded.elements<CanonicalSubsets[sum].elements$ then $\triangleright$ Lex order
@@
1:procedure Check($kSubset,target,theOtherKSubsets,theOtherMemoizedSums,currentSplitSum$)
@@
3:       print “Subset from a single split: ” + $kSubset$
@@
5:    if $(currentSplitSum-kSubset.sum)=target$ then
6:       print “Subset found on a single split using a complement subset: ” + $kSubset$
@@
12:          print “(A+B) subset sum found using both splits: ” + $composedSubset$
@@
15:          print “(A+B’) subset sum found using both splits: ” + $composedSubset$
@@
4:$schedulingQueues\leftarrow\{\{Perm(0,\{\},0),\,KPerm(0,\{\},1)\},\,\{\},\,...,\,\{\}\}$
5:$lookAhead\leftarrow n/16$triangleright Empirically determined
@@
8:procedure ExpandOrSchedule($kPerm,\,splitElements,\,splitMemoizedSums,\,currentCycle$)
@@
16:             $schedulingQueues[currentCycle+currentStep].Add(kPerm)$triangleright Reschedule
17:             $kPerm.LastDeferralIndex\leftarrow kIndex$triangleright Update continuation index
18:             $mustExpand\leftarrow False$triangleright Drop from current enumeration
19:       $kIndex\leftarrow kIndex+1$triangleright Next element in look-ahead block
20:       $currentStep\leftarrow currentStep+1$triangleright Increment step count
21:    return $mustExpand$
@@
24:procedure Solver($S,\,target$)
@@
29:          $Input\leftarrow futureIterations[currentIteration-1]$triangleright Use the schedule array
30:       for  $i\leftarrow 0$ to $(n/4)-1$  do $\triangleright$ For the required columns in the split

### v3

@@ Abstract.
jesus.salas@gmail
@@ 3.2. Unique Subset Sums
For clarity, we assume throughout that $S\subset\mathbb{Z}_{\geq 0}$. The methods generalize naturally to inputs containing 0 (introducing a trivial collision) or negative values (handled by symmetric memoization). In the worst case, storing all unique subset sums requires $O^{*}(2^{n/2})$ space, though this cost is typically much lower in collision-rich instances.
@@ 4. Prior Work
The Subset Sum problem has attracted extensive research due to its NP-completeness and practical significance. Early approaches, such as Bellman’s dynamic programming algorithm [16], offered pseudopolynomial-time solutions with time complexity O(nt)O(n\,t) (where $t$ is the target sum). Over time, several refined DP-based techniques have been proposed to mitigate this cost when $t$ remains large. For instance, Koiliaris and Xu [17] introduced a faster pseudopolynomial-time algorithm running in $\widetilde{O}(t\sqrt{n})$ by leveraging FFT-based convolutions for partial-sum computations. More recently, Bringmann [18] developed a near-linear pseudopolynomial algorithm for Subset Sum under specific parameter regimes, achieving further speedups in cases where the numeric range is not excessively large. Nevertheless, once $t$ grows beyond a moderate threshold, these DP-based methods become infeasible in practice because their complexity remains tied to the numerical size of the target.
@@ 5. Unique Subset Sums Enumerator
Initial step: Start with an INPUT list containing the empty $k$-permutation, $\{\}$.
@@
Once all entries in INPUT have been processed, set $\textbf{INPUT}\leftarrow\textbf{OUTPUT}$ and repeat until no new $k$-permutations are generated.
@@
This strategy eliminates redundant enumeration and significantly reduces both computational overhead and memory usage, while ensuring that all unique subset sums—i.e., all elements of $\Sigma(S)$ (and hence the effective parameter UU)—are captured.
@@ 5.6. Complement Trick for Half-Enumeration
Let $\ell$ be one half of the input, containing $n/2$ elements. Suppose the enumerator enumerates only subsets of size $\leq\lfloor n/4\rfloor$. Let $U_{\max}$ be the number of unique subset sums discovered. Then for any subset $S^{\prime}\subseteq\ell$ of size >n/4>n/4, its sum can be reconstructed via complement logic using previously enumerated subsets of size $\leq n/4$.
@@
Let $S^{\prime}\subseteq\ell$ be a subset of size >n/4>n/4. Then its complement $S^{*}=\ell\setminus S^{\prime}$ has size <n/4<n/4, since $\ell$ has size $n/2$. The enumerator processes all such small subsets and stores their sums. Let $\sigma^{*}=\sum_{x\in S^{*}}x$ be the recorded sum. Then:
@@ 6. Subset-Sum Solver with Double Meet-in-the-Middle Optimization
Let $\sigma(A):=\sum_{a\in A}a$ denote the sum of elements in a subset $A\subseteq S$, in contrast to $\Sigma(A)$, which refers to the full sumset of all subsets of AA
@@ 6.1. Time Complexity Analysis
The enumeration algorithm maintains a dynamic frontier of reachable sums across both halves of the input. Each side’s unique sums are generated via the structure-aware subset sum enumerator described earlier, and combinations are performed only on compatible pairs. For every sum $s$ from the left half, the algorithm checks for a corresponding sum tst-s in the right half, or other lemma-induced matches.
@@ 7. Meet-in-the-Middle Speed-Up via Controlled Aliasing
[…]
@@ 8.1. Overview of the Optimization
Figure 1 shows the number of $k$-subsets generated in each column expansion sub-process for an input instance with $n=48$. In this case, the process is executed $n/4=12$ times, yielding a total of 8 388 6078\,388\,607 $k$-subsets per split, which corresponds to the expected $\frac{2^{n/2}}{2}$ in the worst case. The memoization structure holds only half of the unique subset sums, reflecting the true effective search space.
@@ 8.2. The Look-Ahead and Rescheduling Mechanism
    $\mathsf{futureCycle}=\mathsf{currentCycle}+(\text{position of candidate in block})$
@@ 9.1. Key Factors Influencing Instance Classification and Hardness
Let $n$ denote the number of elements in the set $S$, and let $w=\log_{2}(\max S)$ be the bit-length of the largest element. In addition to the classical measures of additive structure—such as the doubling constant $C$ and additive energy E(S)E(S)—we also consider the extended structural parameters introduced earlier: the linearity factor $\lambda$, the clustering factor $\gamma$, and the duplicate measure $\delta$. These parameters collectively capture the degree of additive redundancy in $S$ and offer a finer-grained understanding of how collisions among subset sums occur.
@@ 10.3. Duplicates Elements
We observe that the most expensive enumeration matches our baseline results, and that each additional duplicate reduces the number of effective unique subset sums by $\penalty 10000\ 25\%$ from the previous run. These sums are tracked using canonical bitmask representations, avoiding redundant credit for distinct but permutationally equivalent subsets.
@@ 12.1. Online Updates as a New Branch
A distinctive benefit of our anytime exploration is that it renders the search process highly governable. Since the algorithm operates in configurable discrete cycles—each producing a set of candidate branches that contribute to UU—we can reassess the search state after every cycle. Once a cycle completes, the scheduler can reorganize the rescheduled branches by prioritizing those likely to yield new unique canonical subset sums, pruning those that are redundant (i.e., those whose sums already exist in $\Sigma(S)$), and even deliberately encouraging collisions to block redundant exploration in unpromising regions of the search space. This adaptive prioritization not only guarantees the delivery of partial solutions at any interruption point but also enables the solver to be controlled and fine-tuned in real time, enhancing overall efficiency.
@@ Appendix A Unique Subset Sums Enumerator Algorithm
1:struct kPerm(sum, elements) { $\triangleright$ Canonical k-permutation
2:  sum $\leftarrow 0$ $\triangleright$ Current subset sum
3:  elements[n] $\leftarrow\{0,0,\ldots\}$ $\triangleright$ Inclusion bitmask
@@
[…]
@@
25:        else if expanded.elements<CanonicalSubsets[sum].elementsexpanded.elements<CanonicalSubsets[sum].elements then $\triangleright$ Lex order
@@
1:procedure Check(kSubset,target,theOtherKSubsets,theOtherMemoizedSums,currentSplitSumkSubset,target,theOtherKSubsets,theOtherMemoizedSums,currentSplitSum)
@@
3:   print “Subset from a single split: ” + kSubsetkSubset
@@
5:  if (currentSplitSumkSubset.sum)=target(currentSplitSum-kSubset.sum)=target then
6:   print “Subset found on a single split using a complement subset: ” + kSubsetkSubset
@@
12:    print “(A+B) subset sum found using both splits: ” + composedSubsetcomposedSubset
@@
15:    print “(A+B’) subset sum found using both splits: ” + composedSubsetcomposedSubset
@@
[…]
@@
8:procedure ExpandOrSchedule(kPerm,splitElements,splitMemoizedSums,currentCyclekPerm,\,splitElements,\,splitMemoizedSums,\,currentCycle)
@@
16:      schedulingQueues[currentCycle+currentStep].Add(kPerm)schedulingQueues[currentCycle+currentStep].Add(kPerm) $\triangleright$ Reschedule
17:      $kPerm.LastDeferralIndex\leftarrow kIndex$ $\triangleright$ Update continuation index
18:      $mustExpand\leftarrow False$ $\triangleright$ Drop from current enumeration
19:   $kIndex\leftarrow kIndex+1$ $\triangleright$ Next element in look-ahead block
20:   $currentStep\leftarrow currentStep+1$ $\triangleright$ Increment step count
21:  return mustExpandmustExpand
@@
24:procedure Solver(S,targetS,\,target)
@@
29:    $Input\leftarrow futureIterations[currentIteration-1]$ $\triangleright$ Use the schedule array
30:   for $i\leftarrow 0$ to (n/4)1(n/4)-1 do $\triangleright$ For the required columns in the split